If x^2+2x=3, then x could equal?
Answer by AnlytcPhil(1806)

If x^2+2x=3, then x could equal?
 
x2 + 2x = 3

Get 0 on the right by subtracting 3 from both sides

x2 + 2x - 3 = 0

Factor the left side

(x+3)(x-1) = 0

One of those factors (x+3) and (x-1) must be = 0 in order for the
right side to be 0. 

Setting x+3 = 0 gives x = -3

Setting x-1 = 0 gives x = 1

So x can be either -3 or 1

Edwin
AnlytcPhil@aol.com